# Complementary Time–Space Tradeoff for Self-Stabilizing Leader Election: Polynomial States Meet Sublinear Time

### v1

@@ 3.3 Cyclic Execution of Sub-protocols
Moreover, for any $r\in[1,n]$, we define $\mathcal{C}_{\mathrm{init}}(\mathsf{D},r)$ as the set of configurations in $\mathcal{C}_{\mathrm{init}}(\mathsf{D})$ where $a.\mathtt{target}=r$ for every agent $a\in A$.
Notation: $R_{\mathrm{name}}=[1,\rho^{2}]\setminus\{r\}$
Variables: $\mathtt{list}\in\{S\subseteq R_{\mathrm{name}}:|S|\leq\rho\}$, $\mathtt{kID}\in R_{\mathrm{name}}\cup\{\bot\}$, $\mathtt{susp}\in\{0,1\}$
[…]
Initially: $a.\mathtt{det}=0$, $a.\mathtt{list}=\emptyset$, $a.\mathtt{kID}=\bot$, $a.\mathtt{susp}=0$
[…]
20       $a_{\mathrm{res}}.\mathtt{kID}\leftarrow a_{\mathrm{ini}}.\mathtt{kID}\leftarrow\begin{cases}a_{\mathrm{res}}.\mathtt{rank}&\textit{if }a_{\mathrm{ini}}.\mathtt{kID}=\bot\\ a_{\mathrm{ini}}.\mathtt{kID}&\textit{otherwise}\end{cases}$
21
[…]
23       $a_{\mathrm{res}}.\mathtt{list}\leftarrow a_{\mathrm{ini}}.\mathtt{list}\cup a_{\mathrm{res}}.\mathtt{list}$
24
[…]
26       $a_{\mathrm{res}}.\mathtt{susp}\leftarrow 1$
27
[…]
30
[…]
Algorithm 2 ${\textsc{Detect}}(r,\rho)$ at an interaction where initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$ meets. Whenever this sub-protocol is invoked, $r=a_{\mathrm{ini}}.\mathtt{target}=a_{\mathrm{res}}.\mathtt{target}$.
3.4 Careful Collision Detection
Given a target rank $r$, the sub-protocol ${\textsc{Detect}}(r,\rho)$, shown in Algorithm 2, detects whether two or more agents share rank $r$, storing the result in each agent’s $\mathtt{det}$ variable. When agents collide on some other rank $r^{\prime}\neq r$, ${\textsc{Detect}}(r,\rho)$ may detect a collision even if only one agent currently holds rank $r$. Such misdetections are harmless: any detected collision should trigger a rank reassignment.
Every agent maintains a set variable $\mathtt{list}\subseteq R_{\mathrm{name}}$ with $|\mathtt{list}|\leq\rho$, where $R_{\mathrm{name}}=[1,\rho^{2}]\setminus\{r\}$, a variable $\mathtt{kID}\in R_{\mathrm{name}}\cup\{\bot\}$, and a Boolean variable $\mathtt{susp}\in\{0,1\}$. When an agent $a$ switches to mode $\mathsf{D}$ (line 15), it initializes $a.\mathtt{det}=0$, $a.\mathtt{list}=\emptyset$, $a.\mathtt{kID}=\bot$, and $a.\mathtt{susp}=0$. Once any agent sets its $\mathtt{det}$ flag to $1$, this information propagates to the entire population via the epidemic protocol (line 27). We now explain how ${\textsc{Detect}}(r,\rho)$ ensures that, whenever a collision exists, at least one agent raises its $\mathtt{det}$ flag within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$.
We classify agents into four roles: kings, vassals, ronins, and commoners. An agent of rank $r$ is a king. A non-king agent $a$ with $a.\mathtt{rank}\in R_{\mathrm{name}}$ is a vassal if $a.\mathtt{rank}\in a.\mathtt{list}$, and a ronin (i.e., a masterless warrior) otherwise. All other agents (those with rank in $[\,\rho^{2}+1,n]\setminus\{r\}$) are commoners. This protocol concerns only kings, vassals, and ronins; commoners are ignored except by direct collision detection: whenever two agents of the same rank interact, the responder immediately sets $\mathtt{det}=1$ (line 27). Direct collision detection simplifies our analysis: we may, without loss of generality, restrict attention to initial configurations with few colliding pairs, particularly those with few kings and many ronins, thanks to the following simple observation.
Observation 1 (Generalization of Lemma 3 of [8]).
In a population of $n$ agents with $x\geq 1$ colliding pairs (i.e., unordered pairs of agents sharing the same rank), the direct collision detection protocol detects a collision within $O(n^{2}/x)$ interactions both in expectation and with probability $\Omega(1)$.
Proof.
Each interaction involves a colliding pair with probability at least $x/\binom{n}{2}=\Omega(x/n^{2})$, so the expected detection time is $O(n^{2}/x)$. The constant-probability bound then follows by Markov’s inequality. ∎
The goal of a king is to detect the presence of any other king. Roughly, each king adopts up to $\rho$ vassals and records their ranks in its $\mathtt{list}$. If a vassal encounters another king, it sets its $\mathtt{susp}$ flag to $1$, and this flag propagates to all of the king’s vassals via a variant of the epidemic protocol. When the king later interacts with a trusted vassal whose $\mathtt{susp}=1$, it sets its own $\mathtt{det}$ flag to $1$. This high-level overview omits many rules needed to achieve fast detection and avoid false positives. In the remainder of this subsection, we formalize this process and prove it satisfies both:
  • Liveness (Lemma 5): for any $C\in\mathcal{C}_{\mathrm{init}}(\mathsf{D},r)$ with at least two agents of rank $r$, a collision is detected with probability $\Omega(1)$ within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions; and
  • Safety (Lemma 6): once the population under $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ reaches a safe configuration, no false positives occur.
Let $a_{K}$ be a king. When $a_{K}$ (as initiator) meets a ronin $b$, $a_{K}$ adopts $b$ as a vassal: it adds $b.\mathtt{rank}$ to $a_{K}.\mathtt{list}$, and $b$ copies this into $b.\mathtt{list}$ (line 19). If $a_{K}.\mathtt{list}$ already contains $\rho$ ranks or includes $b.\mathtt{rank}$, then $a_{K}$ ignores $b$. Simultaneously, $a_{K}$ records the rank of its first vassal in $a_{K}.\mathtt{kID}$ as its kingdom identifier and disseminates it to all adopted vassals (line 20). As we will show later, we may assume w.l.o.g. that no other king shares the same kingdom identifier as $a_{K}$. We say that an agent $a$ is saturated if $|a.\mathtt{list}|=\rho$, while an agent $a$ trusts an agent $b$ if $a.\mathtt{kID}=b.\mathtt{kID}$, $a.\mathtt{list}\subseteq b.\mathtt{list}$, and $b$ is saturated. Vassals always attempt to acquire the complete vassal list from any trusted agent, especially a king: whenever a vassal $a_{\mathrm{res}}$ (as responder) meets a trusted agent $a_{\mathrm{ini}}$, the vassal $a_{\mathrm{res}}$ adds all ranks in $a_{\mathrm{ini}}.\mathtt{list}$ to its own list if capacity allows (lines 21–22). If a vassal $a_{\mathrm{res}}$ interacts with a king with a different kingdom identifier, then $a_{\mathrm{res}}$ sets its $\mathtt{susp}$ flag to $1$, indicating a suspected collision (line 23). Note that this event can occur even when there is exactly one king, if $a_{\mathrm{res}}$ is outdated—i.e., the unique king $a_{\mathrm{ini}}$ has already started the next cycle and obtains a new kingdom identifier. Although the phase clock synchronizes the population w.h.p., making this event exceedingly rare, we must handle it to solve SS-RK (whereas LS-RK could safely ignore it). After a vassal raises its $\mathtt{susp}$ flag, this information spreads among vassals via a variant of the epidemic protocol: the raised $\mathtt{susp}$ flag propagates from an initiator to a responder if and only if they trust each other (line 23). Once a king’s $\mathtt{susp}$ flag is set, the king finally sets its $\mathtt{det}$ flag.
Lemma 5.
Starting from any configuration $C\in\mathcal{C}_{\mathrm{init}}(\mathsf{D},r)$ with $r\in[1,n]$, in which at least two agents hold the target rank $r$, the population under $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ reaches a configuration where every agent has $\mathtt{det}=1$ within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$.
Proof.
Fix any king $a_{K}$ in $C$. It suffices to show that $a_{K}$ raises its $\mathtt{det}$ flag within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$; by Lemma 2, this flag then propagates to all agents within $O(n\log n)=O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p. We may assume w.l.o.g. that $\rho=\omega(1)$, since if $\rho=O(1)$, direct collision detection alone raises $\mathtt{det}$ within $O(n^{2})=O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$, as desired.
Let $x_{C}$, $n_{K}$, and $n_{R}$ be the numbers of colliding pairs (i.e., unordered pairs of agents sharing the same rank), kings, and ronins, respectively, in configuration $C$. By Observation 1, we may assume w.l.o.g. that $x_{C}=o(\rho/\log\rho)$. Since $x_{C}\geq\binom{n_{K}}{2}$, this implies $n_{K}=o(\sqrt{\rho/\log\rho})$. Because $C\in\mathcal{C}_{\mathrm{init}}(\mathsf{D},r)$ contains no vassals, only ronins occupy ranks in $R_{\mathrm{name}}=[1,\rho^{2}]\setminus\{r\}$. If $n_{R}<|R_{\mathrm{name}}|$, the pigeonhole principle gives $x_{C}\geq|R_{\mathrm{name}}|-n_{R}$. Hence $n_{R}=|R_{\mathrm{name}}|-o(\rho/\log\rho)=\Omega(\rho^{2})$.
We may assume w.l.o.g. that $a_{K}$ obtains a unique kingdom identifier. Call an agent colliding if it shares its rank with another agent; at most $2x_{C}$ agents are colliding in $C$. Since each of the $n_{K}$ kings adopts at most $\rho$ vassals, there are always at least $n_{R}-n_{K}\cdot\rho=\Omega(\rho^{2})$ ronins. Hence, when $a_{K}$ meets a ronin for the first time, that ronin is non-colliding with probability $1-2x_{C}/(n_{R}-n_{K}\cdot\rho)=1-o(1)$. Therefore, with probability $1-o(1)$, $a_{K}$’s first-vassal rank is unique, giving it a unique kingdom identifier. This negligible failure probability does not affect the lemma.
By Lemma 3, we may lengthen each phase by an arbitrarily large constant factor $c_{\mathrm{min}}$. Hence, by Markov’s inequality, it suffices to show that $a_{K}$ raises its $\mathtt{det}$ flag within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ expected interactions, assuming no agent leaves mode $\mathsf{D}$.
Since $\Omega(\rho^{2})$ ronins always exist as mentioned above, $a_{K}$ meets a ronin with probability $\Omega(\rho^{2}/n^{2})$ per step. Therefore, $a_{K}$ collects $\rho$ distinct ranks (allowing for up to $x_{C}$ redundant collisions) and becomes saturated within $O((\rho+x_{C})\,n^{2}/\rho^{2})=O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ expected interactions. Let $V=\{v_{1},\dots,v_{\rho}\}$ be the vassals of $a_{K}$, and let $R=\{v.\mathtt{rank}:v\in V\}$ be their ranks. Since $a_{K}$ and every $v\in V$ share the unique identifier $g=a_{K}.\mathtt{kID}$, no vassal trusts agents outside $V\cup\{a_{K}\}$, so each $v\in V$ only ever learns ranks in $R$. Therefore, each $v\in V$ always trusts $a_{K}$ and any saturated vassal in $V$ (after $a_{K}$ becomes saturated), reducing the spread of saturation to an epidemic on $V\cup\{a_{K}\}$: a vassal in $V$ becomes saturated each time it (as responder) meets any saturated agent in $V\cup\{a_{K}\}$. Interactions within $V\cup\{a_{K}\}$ occur with probability $\binom{\rho+1}{2}/\binom{n}{2}=\Theta(\rho^{2}/n^{2})$ per step. Since the epidemic on the entire population completes within $O(n\log n)$ interactions w.h.p. (hence also in expectation), this restricted epidemic disseminates $R$ to all vassals in $V$ within $O(\nicefrac{{n^{2}}}{{\rho^{2}}}\cdot\rho\log\rho)=O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ expected interactions.
Subsequently, within $O(n^{2}/\rho)$ expected interactions, some vassal in $V$ interacts with a king other than $a_{K}$, which must have a different kingdom identifier, causing that vassal to set its $\mathtt{susp}$ flag. Since all saturated vassals in $V$ trust one another, this $\mathtt{susp}$ flag then spreads to all agents in $V\cup\{a_{K}\}$ via the epidemic on $V\cup\{a_{K}\}$ within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ expected interactions. During this spread, $a_{K}$ meets a vassal with $\mathtt{susp}=1$ and raises its own $\mathtt{det}$ flag, as desired. ∎
Definition 5 (Safe Configuration).
Let $\mathcal{S}$ be the set of configurations of $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ satisfying: (i) all agents have distinct ranks; (ii) no agent is in the mode $\mathsf{R}$; (iii) every agent has $\mathtt{det}=0$; (iv) every king and ronin has $\mathtt{susp}=0$, i.e., for all $a\in A$,
[…]
(v) If a king’s $\mathtt{list}$ contains a vassal’s $\mathtt{list}$, then the vassal has $\mathtt{susp}=0$ and shares the same kingdom identifier as the king, i.e., for all $a,b\in A$,
[…]
Lemma 6.
$\mathcal{S}$ is closed under interactions: any configuration in $\mathcal{S}$ remains in $\mathcal{S}$ after a single interaction.
Proof.
Let $C\in\mathcal{S}$, and suppose $C$ transitions to $C^{\prime}$ by one interaction between initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$. We show $C^{\prime}\in\mathcal{S}$.
First, $C^{\prime}$ clearly satisfies (i) and (ii): no rank changes (since $a_{\mathrm{res}}\notin\mathsf{R}$) and no agent switches to $\mathsf{R}$ (since $a_{\mathrm{res}}.\mathtt{det}=0$). It remains to check (iii), (iv), and (v). If any of these were violated in $C^{\prime}$, then either $a_{\mathrm{ini}}$ or $a_{\mathrm{res}}$ must have updated one of the variables $\mathtt{det},\mathtt{susp},\mathtt{mode},\mathtt{target},\mathtt{list},\mathtt{kID}$. There are two cases:
Case (A): $a_{\mathrm{res}}$ switches its mode to $\mathsf{D}$.
Then $a_{\mathrm{res}}$ reinitializes $\mathtt{det}=0$, $\mathtt{list}=\emptyset$, $\mathtt{kID}=\bot$, and $\mathtt{susp}=0$. This preserves (iii) and (iv). For (v), note $a_{\mathrm{res}}$ is not a vassal in $C^{\prime}$, and if $a_{\mathrm{res}}$ is a king, its empty $\mathtt{list}$ cannot contain any vassal’s list, which must not be empty. Thus (v) holds.
Case (B): $a_{\mathrm{ini}}$ and $a_{\mathrm{res}}$ share the same target rank $r$ and execute ${\textsc{Detect}}(r,\rho)$.
By (i), exactly one agent—call it $a_{K}$—has rank $r$. If $a_{K}.\mathtt{mode}\neq\mathsf{D}$ or $a_{K}.\mathtt{target}\neq r$, then (iii)–(v) remain intact. So assume $a_{K}.\mathtt{mode}=\mathsf{D}$ and $a_{K}.\mathtt{target}=r$.
Subcase (B–1): $a_{\mathrm{res}}$ is a vassal in $C$.
Then $a_{\mathrm{res}}$ cannot violate (iii) or (iv). To violate (v), $a_{\mathrm{res}}$ would have to raise $\mathtt{susp}=1$ while still having $a_{\mathrm{res}}.\mathtt{list}\subseteq a_{K}.\mathtt{list}$ (no vassal changes its $\mathtt{kID}$). That requires the initiator to be a trusted agent: if $a_{\mathrm{ini}}=a_{K}$, then by (v) in $C$, $a_{\mathrm{res}}$ has $\mathtt{susp}=0$; if $a_{\mathrm{ini}}$ is another vassal, line 24 raises $\mathtt{susp}$ only when $a_{\mathrm{ini}}.\mathtt{list}=a_{\mathrm{res}}.\mathtt{list}$ and $a_{\mathrm{ini}}.\mathtt{susp}=1$, which violates (v). Hence no violation occurs.
Subcase (B–2): $a_{\mathrm{res}}=a_{K}$.
Then only lines 25–27 can update $\mathtt{susp}$. But since all vassals trusted by $a_{K}$ have $\mathtt{susp}=0$ in $C$, $a_{K}$ cannot set $\mathtt{susp}=1$. Thus (iii) and (iv) are preserved. Since a king never updates its $\mathtt{list}$ when it is a responder, (v) clearly holds.
In every case, $C^{\prime}\in\mathcal{S}$. ∎
[…]
@@ 3.4 Careful Collision Detection
[…]
@@
39
40
41if  $p_{2}\geq T_{3.5}$ and $p_{2}\equiv a_{\mathrm{res}}.\mathtt{parity}\pmod{2}$  then
@@
49
50if $p_{1}=p_{2}\geq T_{3.5}\land a_{\mathrm{res}}.\mathtt{nonce}<a_{\mathrm{ini}}.\mathtt{nonce}$ then
@@
53
54
Algorithm 3 Rank at an interaction where initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$ meets.
3.5 Sublinear-Time, Polynomial-State Ranking
@@
Within this protocol, $\mathtt{rank}$ ranges over $[1,n]\cup\{\bot\}$ rather than $[1,n]$. If $a.\mathtt{mode}$ leaves $\mathsf{D}$ while $a.\mathtt{rank}=\bot$, we reset $a.\mathtt{rank}$ to some value in $[1,n]$ (e.g. $1$).
The module Rank, shown in Algorithm 3, consists of two parts. In the first part (lines 28–35), it assigns $n-m$ agents ranks in $[1,n-m]$ in parallel. In the second part (lines 36–47), it assigns the remaining $m$ agents ranks in $[n-m+1,n]$ sequentially. The first part requires $O\left(\nicefrac{{n^{2}}}{{m}}\cdot\log n\right)$ interactions w.h.p., and the second requires $O(mn\log n)$ interactions w.h.p. With $m=\sqrt{n}$, the total becomes $O\left(n^{3/2}\log n\right)$, as claimed.
Recall that this subprotocol runs over phases $[T_{3},T_{4}-2]$, where $T_{4}=T_{3}+2c_{T}\left\lceil\sqrt{n}\right\rceil$. To execute its two parts sequentially, we split these phases nearly in half into $[T_{3},T_{3.5}-1]$ and $[T_{3.5},T_{4}-2]$, where
@@
We design the first part, inspired by the $O(n)$-time, $O(n)$-state SS-RK protocol Optimal-Silent-SSR of Burman et al. [14]. Whereas Optimal-Silent-SSR assigns all $n$ ranks, our first part assigns only $n-m$ ranks, deferring the remaining $m$ ranks to the second part. This reduces the interaction complexity from $O(n^{2})$ in expectation (and $O(n^{2}\log n)$ w.h.p.) to $O(n^{3/2}\log n)$ interactions w.h.p.
The first part assigns $n-m$ ranks using a variable $\mathtt{index}\in[0,\lceil\lg n\rceil]$. We call an agent with $\mathtt{rank}=\bot$ a null agent. Each non-null agent $a$ interprets the pair $(a.\mathtt{rank}-1,\,a.\mathtt{index})$ as a $\lceil\lg n\rceil$-bit integer whose first $a.\mathtt{index}$ bits are fixed and whose remaining bits are masked:
@@
where ‘*’ denotes a masked bit. Note that subtracting 1 from $a.\mathtt{rank}$ maps ranks in $[1,n-m]$ to the zero-based range $[0,n-m-1]$. When $a.\mathtt{index}=\lceil\lg n\rceil$, $a.\mathtt{rank}$ is fully determined and no longer changes. Upon switching to mode $\mathsf{R}$, agents use the leader $a_{L}$, maintained by module $\mathcal{P}_{{\textsc{LSLE}}}$ (line 1): the unique leader $a_{L}$ sets $\mathtt{rank}=1$, while all others set $\mathtt{rank}=\bot$. All agents also initialize $\mathtt{index}=0$. Hence initially only $a_{L}$ is non-null with $(\mathtt{rank}-1,\mathtt{index})=(0,0)$, and all others are null. The first part proceeds when a non-null initiator $a_{\mathrm{ini}}$ with $a_{\mathrm{ini}}.\mathtt{index}<\lceil\lg n\rceil$ meets a null responder $a_{\mathrm{res}}$. Then $a_{\mathrm{ini}}.\mathtt{index}$ increments by one—replacing the leftmost masked bit with 00 (line 31)—and $r_{\mathrm{new}}-1$ is obtained by setting that bit to $1$ in the binary representation of $a_{\mathrm{ini}}.\mathtt{rank}-1$ (line 32). If $r_{\mathrm{new}}\leq n-m$, the responder sets $(\mathtt{rank},\mathtt{index})=(r_{\mathrm{new}},a_{\mathrm{ini}}.\mathtt{index})$ (lines 33–35).
In the first part, at most $n-m$ ranks are generated, so at least $m$ null agents always remain. This key property yields faster stabilization. Fix any $r\in[1,n-m]$. At each time step $t$, there is exactly one agent $a_{r}(t)\in A$ satisfying
@@
The second part assigns $m$ ranks sequentially, using the phase clock that $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ invokes outside this module (line 10). In each round of $c_{T}$ phases, all null agents contend to win a new rank. With high probability, exactly one agent wins per round and obtains its assigned rank. This follows a standard leader-election technique in population protocols [27, 31, 22]. Specifically, when an agent first reaches phase $p\in[T_{3.5},T_{4}-2]$—detectable via a binary variable $\mathtt{parity}\in\{0,1\}$—it executes:
  • If $p-T_{3.5}$ is divisible by $c_{T}$, a null agent sets its $\mathtt{cand}$ flag to $1$ (line 42), becoming a candidate for rank $n-m+1+(p-T_{3.5})/c_{T}$. A candidate that retains $\mathtt{cand}=1$ for the next $c_{T}$ phases wins that rank; non-null agents always reset $\mathtt{cand}=0$.
  • Every agent with $\mathtt{cand}=1$ chooses its nonce u.a.r  from $[1,n^{2}]$ (derandomizable—see Note 3); agents with $\mathtt{cand}=0$ set $\mathtt{nonce}=0$.
In each phase, the maximum nonce propagates to the entire population w.h.p. via the epidemic protocol among agents in the same phase (lines 45–47). Any agent that observes a larger nonce resets its $\mathtt{cand}$ flag to 00 (line 46).
Let $n_{\mathrm{cand}}$ be the number of candidates in a given phase. They draw nonces $x_{1},\dots,x_{n_{\mathrm{cand}}}$ u.a.r. from $[1,n^{2}]$. For each $i\in[2,n_{\mathrm{cand}}]$, $\Pr[x_{i}=\max\{x_{1},\dots,x_{i-1}\}]=1/n^{2}$, so by a union bound over $i=2,\dots,n_{\mathrm{cand}}$, there is a unique maximum nonce with probability $1-O(n_{\mathrm{cand}}/n^{2})=1-O(1/n)$. Hence, by choosing $c_{T}=\Theta(1)$ sufficiently large, each round selects exactly one winner w.h.p.
@@
In Rank, each non-null agent generates a random number from $[1,n^{2}]$, making the protocol randomized. However, we can derandomize it by exploiting the scheduler’s randomness: whenever agent $a$ is chosen for an interaction, it becomes initiator with probability $1/2$ (and responder otherwise), yielding one random bit per interaction. By Lemma 3, each phase $i$ involves $\Theta(n\log n)$ interactions w.h.p., of which $a$ participates in $\Theta(\log n)$ w.h.p. (for sufficiently large constant $c_{\mathrm{min}}$). Thus $a$ obtains $\Theta(\log n)$ random bits per phase—enough to generate a number in $[1,n^{2}]$ for the next phase. This assumes $n^{2}$ is a power of 2, but replacing $n^{2}$ with $2^{\lceil 2\lg n\rceil}$ only increase the success probability. Note that this derandomization does not make nonces independent across agents. In Rank, such dependence only accelerates the elimination of losing candidates. If full independence is required, one can use the derandomization method of Burman [14] (see Section 6 of the arXiv version [15]).
3.6 Finding a Target

### v2

@@ 6 Cyclic Execution of Sub-protocols
Moreover, for any $r\in[1,n]$, let $\mathcal{C}_{\mathrm{init}}(\mathsf{D},r)$ denote the set of configurations in $\mathcal{C}_{\mathrm{init}}(\mathsf{D})$ in which $a.\mathtt{target}=r$ for every agent $a\in A$.
Notation: $m=\lceil\sqrt{n}\rceil$, $\mathit{Phase}(a)=\lfloor a.\mathtt{clock}/c_{M}\rfloor$
@@
Let $p_{1}=\mathit{Phase}(a_{\mathrm{ini}}),\ p_{2}=\mathit{Phase}(a_{\mathrm{res}})$
@@
35
36
37if $p_{2}\geq T_{3.5}$ and $p_{2}\equiv a_{\mathrm{res}}.\mathtt{parity}\pmod{2}$ then
@@
45
46if $p_{1}=p_{2}\geq T_{3.5}\land a_{\mathrm{res}}.\mathtt{nonce}<a_{\mathrm{ini}}.\mathtt{nonce}$ then
@@
49
Algorithm 3 Rank at an interaction where initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$ meet.
7 Sublinear-Time, Polynomial-State Ranking
@@
Within this protocol, $\mathtt{rank}$ ranges over $1,n]\cup\{\bot\}$ rather than $[1,n]$. If $a.\mathtt{mode}$ leaves $\mathsf{R}$ while $a.\mathtt{rank}=\bot$, we reset $a.\mathtt{rank}$ to some value in $[1,n]$ (e.g. $1$).
The module Rank, shown in Algorithm 3, consists of two parts. In the first part (lines 25–32), it assigns nmn-m agents ranks in [1,nm][1,n-m] in parallel. In the second part (lines 33–44), it assigns the remaining $m$ agents ranks in [nm+1,n][n-m+1,n] sequentially. The first part requires $O\left(\nicefrac{{n^{2}}}{{m}}\cdot\log n\right)$ interactions w.h.p., and the second requires $O(mn\log n)$ interactions w.h.p. With $m=\sqrt{n}$, the total becomes $O\left(n^{3/2}\log n\right)$, as claimed.
Recall that this subprotocol runs over phases $T_{3},T_{4}-2]$, where $T_{4}=T_{3}+2c_{T}\left\lceil\sqrt{n}\right\rceil$. To execute its two parts sequentially, we split these phases nearly in half into $T_{3},T_{3.5}-1]$ and $T_{3.5},T_{4}-2]$, where
@@
We design the first part, inspired by the $O(n)$-time, $O(n)$-state SS-RK protocol Optimal-Silent-SSR of [BCC+21a]. Whereas Optimal-Silent-SSR assigns all $n$ ranks, our first part assigns only nmn-m ranks, deferring the remaining $m$ ranks to the second part. This reduces the interaction complexity from $O(n^{2})$ in expectation (and $O(n^{2}\log n)$ w.h.p.) to $O(n^{3/2}\log n)$ interactions w.h.p.
The first part assigns nmn-m ranks using a variable $\mathtt{index}\in[0,\lceil\lg n\rceil]$. We call an agent with $\mathtt{rank}=\bot$ a null agent. Each non-null agent $a$ interprets the pair $(a.\mathtt{rank}-1,\,a.\mathtt{index})$ as a $\lceil\lg n\rceil$-bit integer whose first $a.\mathtt{index}$ bits are fixed and whose remaining bits are masked:
@@
where ‘*’ denotes a masked bit. Note that subtracting 1 from $a.\mathtt{rank}$ maps ranks in [1,nm][1,n-m] to the zero‐based range [0,nm1][0,n-m-1]. When $a.\mathtt{index}=\lceil\lg n\rceil$, $a.\mathtt{rank}$ is fully determined and no longer changes. Upon switching to mode $\mathsf{R}$, agents use the leader $a_{L}$, maintained by module $\mathcal{P}_{{\textsc{LSLE}}}$ (line 8): the unique leader $a_{L}$ sets $\mathtt{rank}=1$, while all others set $\mathtt{rank}=\bot$. All agents also initialize $\mathtt{index}=0$. Hence initially only $a_{L}$ is non-null with $(\mathtt{rank}-1,\mathtt{index})=(0,0)$, and all others are null. The first part proceeds when a non‐null initiator $a_{\mathrm{ini}}$ with $a_{\mathrm{ini}}.\mathtt{index}<\lceil\lg n\rceil$ meets a null responder $a_{\mathrm{res}}$. Then $a_{\mathrm{ini}}.\mathtt{index}$ increments by one (line 28)—replacing the leftmost masked bit with 0—and $r_{\mathrm{new}}-1$ is obtained by setting that bit to $1$ in the binary representation of $a_{\mathrm{ini}}.\mathtt{rank}-1$ (line 29). If $r_{\mathrm{new}}\leq n-m$, the responder sets $(\mathtt{rank},\mathtt{index})=(r_{\mathrm{new}},a_{\mathrm{ini}}.\mathtt{index})$ (lines 30–32).
In the first part, at most nmn-m ranks are generated, so at least $m$ null agents always remain. This key property yields faster stabilization. Fix any $r\in[1,n-m]$. At each time step $t$, there is exactly one agent $a_{r}(t)\in A$ satisfying
@@
The second part assigns $m$ ranks sequentially, using the phase clock that $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ invokes outside this module (line 17). In each round of $c_{T}$ phases, all null agents contend to win a new rank. With high probability, exactly one agent wins per round and obtains its assigned rank. This follows a standard leader‐election technique in population protocols [MST18, SOI+20, GJLL25]. Specifically, when an agent first reaches phase $p\in[T_{3.5},T_{4}-2]$—detectable via a binary variable $\mathtt{parity}\in\{0,1\}$—it executes:
  • If $p-T_{3.5}$ is divisible by $c_{T}$, a null agent sets its $\mathtt{cand}$ flag to $1$ (line 39), becoming a candidate for rank $n-m+1+(p-T_{3.5})/c_{T}$. A candidate that retains $\mathtt{cand}=1$ for the next $c_{T}$ phases wins that rank; non‐null agents always reset $\mathtt{cand}=0$.
  • Every agent with $\mathtt{cand}=1$ chooses its nonce u.a.r. from $1,n^{2}]$ (derandomizable—see Note 3); agents with $\mathtt{cand}=0$ set $\mathtt{nonce}=0$.
In each phase, the maximum nonce propagates to the entire population w.h.p. via the epidemic protocol among agents in the same phase (lines 42–44). Any agent that observes a larger nonce resets its $\mathtt{cand}$ flag to 0 (line 43).
Let $n_{\mathrm{cand}}$ be the number of candidates in a given phase. They draw nonces $x_{1},\dots,x_{n_{\mathrm{cand}}}$ u.a.r. from $1,n^{2}]$. For each $i\in[2,n_{\mathrm{cand}}]$, $\Pr[x_{i}=\max\{x_{1},\dots,x_{i-1}\}]=1/n^{2}$, so by a union bound over $i=2,\dots,n_{\mathrm{cand}}$, there is a unique maximum nonce with probability $1-O(n_{\mathrm{cand}}/n^{2})=1-O(1/n)$. Hence, by choosing $c_{T}=\Theta(1)$ sufficiently large, each round selects exactly one winner w.h.p.
@@
In Rank, each non-null agent generates a random number from $1,n^{2}]$, making the protocol randomized. However, we can derandomize it by exploiting the scheduler’s randomness: whenever agent $a$ is chosen for an interaction, it becomes initiator with probability $1/2$ (and responder otherwise), yielding one random bit per interaction. By Lemma 6, each phase $i$ involves $\Theta(n\log n)$ interactions w.h.p., of which $a$ participates in $\Theta(\log n)$ interactions w.h.p. (for sufficiently large constant $c_{\mathrm{min}}$). Thus $a$ obtains $\Theta(\log n)$ random bits per phase—enough to generate a number in $1,n^{2}]$ for the next phase. This assumes $n^{2}$ is a power of 2, but replacing $n^{2}$ with $2^{\lceil 2\lg n\rceil}$ only increases the success probability. Note that this derandomization requires a multiplicative factor of $n^{2}$ in the number of states, although this does not affect our results: it suffices for Rank to use only polynomially many states. Moreover, if one wishes to eliminate these $n^{2}$ states, a special phase can be periodically inserted that is dedicated to generating random bits, where agents reuse the variable $\mathtt{nonce}$ to produce $2\log n$ random bits. Note also that this derandomization does not make nonces independent across agents. In Rank, such dependence only accelerates the elimination of losing candidates. If full independence is required, one can use the derandomization method of [BCC+21a] (see Section 6 of the arXiv version [BCC+21b]).
8 Finding a Target